# atomic orbital API Application Programmer Interface

### v1

@@
Evgueni Dinvay Rasmus Vikhamar-Sandberg  and  Luca Frediani evgueni.dinvay@uit.no Department of Chemistry
@@ Abstract.
The multiconfiguration self-consistent-field (MCSCF) method is revisited with the specific focus on the two electron systems for simplicity. A wave function is represented as a linear combination of Slater determinants. Both orbitals and coefficients of this configuration interaction expansion are optimized following the variational principle making use of the Newton optimization technique. It reduces the MCSCF problem to solving a particular differential Newton system, which can be discretized with multiwavelets and solved iteratively.
@@
Schrödinger equation, energy optimization, MCSCF, multiwavelets.
@@ 1. Introduction
We present the first implementation of the Newton optimization for the multiconfiguration self consistent field (MCSCF) method within the framework of multiresolution analysis (MRA) using multiwavelets. Multiconfigurational treatments are required to capture correlation effects. At the same time, it is well known that optimizing MCSCF wavefunctions is challenging and it is widely recognized that the most effective method is a Newton optimization scheme. In traditional methods, this is achieved by expressing the energy in terms of the configurational expansion coefficients and the basis-set coefficients and then computing the second derivatives. Within a multiwavelet framework one operates in a virtually infinite basis set and the approach must therefore be revised, by introducing a Lagrangian and by solving the Newton equations for the Lagrangian gradient. For this prototype study we limit ourselves to the simple case of two-electron closed-shell system, but the method can in principle be extended to an arbitrary number of electrons and any spin symmetry. We show that this method effectively allows to perform MCSCF calculations of both ground and excited states within the framework of MRA.
@@
Here, we note that this adaptive discretization technique is particularly useful in computational chemistry, where it effectively handles the singular Coulomb interaction between an electron and the nucleus, as well as the resulting cusp in orbitals [9, 11, 13, 21]. Moreover, its success in chemistry applications is largely due to the approximation of the resolvent:
$H(\mu)=\left(-\Delta+\mu^{2}\right)^{-1}\approx\sum_{j}c_{j}e^{t_{j}\Delta}$ (1.1)
@@
based on the heat semigroup [10]. Multiresolution quantum chemistry is rapidly evolving, and we refer readers to a comprehensive review for a broader perspective [5]. An overview of the latest advancements will also appear in an upcoming paper [20].
We use the Lagrangian formalism for the constrain minimization of MCSCF. This is rather uncommon in the field, since all the stationary points of the Lagrangian $\mathcal{L}$ are saddle points. On the other hand, it allows us to introduce the Green’s function in the problem naturally. As a result, one can rewrite the differential Newton system in the integral form.
@@ 2. Notations
Throughout the text we use the following notations $\varphi,\varphi_{i},\delta\varphi_{i}$ for functions from $L^{2}\left(\mathbb{R}^{3}\right)$, that are referred to as orbitals. In this paper the focus is on closed shell two electron systems. Therefore, a Slater determinant associated to a spatial orbital $\varphi_{m}$ is denoted by $\left|m\overline{m}\right\rangle,$ implying spin up and down. The single electron energy operator is denoted by
@@
The following standard [19] notations
[…]
and
[…]
are extensively in use. Very often in the text one encounters a convolution of the form
$J(i\delta k)=\frac{1}{|x|}*(\varphi_{i}\delta\varphi_{k})$
@@
that is a function in $L^{2}\left(\mathbb{R}^{3}\right)$. In computational chemistry literature, it is normally associated with either Coulomb or exchange operators, which justifies the choice of a such notation. Finally, we write shortly
$R_{i}=\left(-\frac{c_{i}^{2}}{2}\Delta-\varepsilon_{ii}\right)^{-1}=\frac{2}{c_{i}^{2}}H\left(\sqrt{-\frac{2\varepsilon_{ii}}{c_{i}^{2}}}\right),$
where the so called orbital energies $\varepsilon_{ii}$ are negative. These numbers $\varepsilon_{ii}$ will be always assumed negative, and so lying in the resolvent set of the kinetic energy operator. During numerical simulations it may happen at some iterations, that this condition is violated. In that case it would mean that the iterative procedure is far from convergence, and that it is safe to flip the sign by multiplying the corresponding incorrect orbital energies $\varepsilon_{ii}$ by $-1$. As it will be clear below, this is in line with the level shift of Hessian in Newton’s method.
3. Newton scheme for SCF of a two electron system
Here we consider a simpler problem of application of Newton optimization for search of Hartree-Fock solution. It turns out to be a saddle point in the Lagrangian formalism, of course. We want to see if (shifted) Newton iterations converge to the Hartree-Fock solution.
The total energy
$E(\varphi)=2(1|h|1)+(11|11)$
Introduce the Lagrangian
$\mathcal{L}=E-2\varepsilon\left(\left\|\varphi\right\|^{2}-1\right)$
and consider
$\left\langle\,\nabla\mathcal{L}(\varphi,\varepsilon),(\delta\varphi,\delta\varepsilon)\,\right\rangle=d\mathcal{L}(\varphi,\varepsilon)(\delta\varphi,\delta\varepsilon)=dE(\varphi)(\delta\varphi)-4\varepsilon\int\varphi\delta\varphi-2\delta\varepsilon\left(\left\|\varphi\right\|^{2}-1\right)$
Therefore, the gradient has the coordinates
$\frac{\delta\mathcal{L}}{\delta\varphi}$ $=4\left(h+J(11)-\varepsilon\right)\varphi$ (3.1)
$\frac{\partial\mathcal{L}}{\partial\varepsilon}$ $=-2\left(\left\|\varphi\right\|^{2}-1\right)$
The shifted Newton equation
$(d\nabla\mathcal{L}(\varphi,\varepsilon)+4\lambda)(\delta\varphi,\delta\varepsilon)=-\nabla\mathcal{L}(\varphi,\varepsilon),\quad\lambda\geqslant 0,$
where
$\frac{1}{4}d\nabla\mathcal{L}(\varphi,\varepsilon)(\delta\varphi,\delta\varepsilon)=\begin{pmatrix}h\delta\varphi-\varepsilon\delta\varphi+J(11)\delta\varphi+2J(1\delta 1)\varphi-\delta\varepsilon\varphi\\ -\int\varphi\delta\varphi\end{pmatrix}.$
The self consistent form consists of the following two equations
$\begin{pmatrix}\left\|\varphi\right\|^{2}&-\lambda\\ \lambda&-1\end{pmatrix}\begin{pmatrix}\delta\varepsilon\\ \int\varphi\delta\varphi\end{pmatrix}=\begin{pmatrix}(\delta 1|h-\varepsilon|1)+3(\delta 11|11)+(1|h-\varepsilon|1)+(11|11)\\ \frac{1}{2}\left(\left\|\varphi\right\|^{2}-1\right)\end{pmatrix}$ (Energy)
or
$\begin{pmatrix}\left\|\varphi\right\|^{2}&\varepsilon-\lambda\\ \lambda&-1\end{pmatrix}\begin{pmatrix}\delta\varepsilon\\ \int\varphi\delta\varphi\end{pmatrix}=\begin{pmatrix}(\delta 1|h|1)+3(\delta 11|11)+(1|h-\varepsilon|1)+(11|11)\\ \frac{1}{2}\left(\left\|\varphi\right\|^{2}-1\right)\end{pmatrix}$ (Energy)
and
$\delta\varphi=-\varphi+\left(-\frac{\Delta}{2}-\varepsilon+\lambda\right)^{-1}\left((\delta\varepsilon+\lambda)\varphi-\left(V_{\text{nuc}}+J(11)\right)\left(\delta\varphi+\varphi\right)-2J(1\delta 1)\varphi\right)$
that we can solve iteratively. In fact, a direct use of these two equations leads to very slow convergence. One can speed it up slightly by using DIIS as an inner loop. However, to make it faster than other methods, one has to normalize $\varphi$ after each update.
Shift $\lambda>0$ with every step normalization, spoils the convergence as follows
  • $\lambda\ll 1$ affecting insignificantly;
  • $\lambda\sim 1$ leading to divergence;
  • $\lambda\gg 1$ slowly convergence to a poor result;
Without normalization: only $\lambda\ll 1$ preserve convergence. It is not surprising, since in the case of a single determinant with two closed-shell electrons, we have exactly one saddle point and no extrema for Lagrangian.
Finally, we remark
$\begin{pmatrix}\left\|\varphi\right\|^{2}&\varepsilon\\ 0&-1\end{pmatrix}\begin{pmatrix}\delta\varepsilon\\ \int\varphi\delta\varphi\end{pmatrix}=\begin{pmatrix}(\delta 1|h|1)+3(\delta 11|11)+(1|h-\varepsilon|1)+(11|11)\\ \frac{1}{2}\left(\left\|\varphi\right\|^{2}-1\right)\end{pmatrix}$ (Energy)
and
$\delta\varphi=-\varphi+\left(-\frac{\Delta}{2}-\varepsilon\right)^{-1}\left(\delta\varepsilon\varphi-\left(V_{\text{nuc}}+J(11)\right)\left(\delta\varphi+\varphi\right)-2J(1\delta 1)\varphi\right)$
If each time after finding a solution to Newton’s equation we normalize the orbital, then $\int\varphi\delta\varphi=0$ and
$\delta\varepsilon=(\delta 1|h|1)+3(\delta 11|11)+(1|h|1)+(11|11)-\varepsilon.$
Moreover, $\varepsilon$ can be reset before running Newton’s inner loop as $\varepsilon=(1|h|1)+(11|11),$ which is in line with the Hartree-Fock equation. Then the Newton’s system simplifies to
$\delta\varepsilon$ $=(\delta 1|h|1)+3(\delta 11|11),$
$\delta\varphi$ $=-\varphi+\left(-\frac{\Delta}{2}-\varepsilon\right)^{-1}\left(\delta\varepsilon\varphi-\left(V_{\text{nuc}}+J(11)\right)\left(\delta\varphi+\varphi\right)-2J(1\delta 1)\varphi\right)$
Thus accepting the result of its first iteration $\delta\varphi=0\mapsto\delta\varepsilon=0\mapsto\delta\varphi$ as an approximate solution, we obtain
$\varphi+\delta\varphi=-\left(-\frac{\Delta}{2}-\varepsilon\right)^{-1}\left(\left(V_{\text{nuc}}+J(11)\right)\varphi\right),$
that is a very well established simple scheme [11, 9]. This demonstrates usefulness of the Green’s function and explains the success of multiwavelets in computational chemistry.
@@ 4. Two configurations
$\frac{1}{4}\frac{\delta\mathcal{L}}{\delta\varphi_{1}}=c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}$ (4.1)
@@
$c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}=\varepsilon_{11}\varphi_{1}+\varepsilon_{12}\varphi_{2}$ (4.2)
@@
$-\sin 2\theta(2(1|h|1)+(11|11))+\sin 2\theta(2(2|h|2)+(22|22))+2\cos 2\theta(12|12)=0$
@@
acting in $L^{2}\left(\mathbb{R}^{3}\right)^{2}\times\mathbb{R}^{4}.$ Now for a given $w=w_{n}$ we solve the Newton’s equation
$d\mathcal{R}(w)\delta w=-\mathcal{R}(w)$
with respect $\delta w$. This is a linear equation that we solve iteratively. Note that in a basis type method one could potentially invert the Jacobian, though for this the basis size should be prohibitively small. Then the update is $w_{n+1}=w_{n}+\delta w$.
$d\mathcal{R}_{i}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{i}(w)\delta\varphi_{1}+\ldots+\frac{\partial\mathcal{R}_{i}}{\partial\varepsilon_{12}}(w)\delta\varepsilon_{12},\quad i=1,2,$
@@
$\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{11}}(w)\delta\varepsilon_{11}+\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{22}}(w)\delta\varepsilon_{22}+\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{12}}(w)\delta\varepsilon_{12}=-\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}$
@@
$\partial_{\varphi_{1}}\mathcal{R}_{3}(w)\delta\varphi_{1}=-4\sin 2\theta((\delta 1|h|1)+(\delta 11|11))+4\cos 2\theta(\delta 12|12)$
@@
$\frac{\partial\mathcal{R}_{3}}{\partial\theta}(w)\delta\theta=(-2\cos 2\theta(2(1|h|1)+(11|11))+2\cos 2\theta(2(2|h|2)+(22|22))-4\sin 2\theta(12|12))\delta\theta$
Finally,
$d\mathcal{R}_{4}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{4}(w)\delta\varphi_{1}=2\int\varphi_{1}\delta\varphi_{1}$
@@
$d\mathcal{R}_{6}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{6}(w)\delta\varphi_{1}+\partial_{\varphi_{2}}\mathcal{R}_{6}(w)\delta\varphi_{2}=\int\delta\varphi_{1}\varphi_{2}+\int\varphi_{1}\delta\varphi_{2}$
4.1. Algorithm
Firstly, $\delta\theta$ is defined as $\delta\theta=\Theta(\delta\varphi_{1},\delta\varphi_{2},w)$ by
$\delta\theta=-\left(\frac{\partial\mathcal{R}_{3}}{\partial\theta}(w)\right)^{-1}\left(\partial_{\varphi_{1}}\mathcal{R}_{3}(w)\delta\varphi_{1}+\partial_{\varphi_{2}}\mathcal{R}_{3}(w)\delta\varphi_{2}+\mathcal{R}_{3}(w)\right)$

### v2

@@
Evgueni Dinvay and Rasmus Vikhamar-Sandberg evgueni.dinvay@gmail.com Department of Chemistry
@@ Abstract.
The multiconfiguration self-consistent field (MCSCF) method is revisited with a specific focus on two-electron systems for simplicity. The wave function is represented as a linear combination of Slater determinants. Both the orbitals and the coefficients of this configuration interaction expansion are optimized according to the variational principle within the Lagrangian formalism, using a Newton optimization scheme. This reduces the MCSCF problem to solving a particular differential Newton system, which can be discretized with multiwavelets and solved iteratively.
@@
Schrödinger equation, energy optimization, MCSCF, multiwavelets
@@ 1. Introduction
We present the first implementation of a Newton optimization scheme for the multiconfiguration self consistent field (MCSCF) method within the framework of multiresolution analysis (MRA) using multiwavelets. Multiconfigurational approaches are essential for capturing correlation effects in quantum chemistry. However, the optimization of MCSCF wavefunctions remains a challenging nonlinear problem due to the coupled dependence on both configurational coefficients and orbitals. It is widely recognized that second-order methods, in particular Newton-type algorithms, provide the most efficient route to convergence when reliable second-derivative information is available.
In conventional formulations, the MCSCF energy is expressed in terms of a finite basis representation, and the Newton step is derived from explicit second derivatives with respect to both configuration interaction coefficients and orbital rotation parameters [10]. In contrast, within a multiwavelet framework, one operates in a virtually infinite basis set, where such parametrizations are neither natural nor numerically advantageous. This necessitates a reformulation of the optimization problem at the functional (basis independent) level.
In this work, we adopt a Lagrangian formulation of the MCSCF problem and derive the corresponding Newton equations directly in function space. This approach enables the incorporation of Green’s function techniques and leads naturally to an integral formulation of the Newton step. The use of resolvent operators plays a central role in this construction and is closely aligned with the multiresolution representation.
For this prototype study we restrict ourselves to the case of a two-electron closed-shell system. While this setting allows for a transparent exposition of the method, the formulation can be generalized and extended to systems with an arbitrary number of electrons and general spin symmetry.
We emphasize that the present formulation is closely related to recent developments in geometric optimization. In particular, the variational parameters of the MCSCF problem naturally reside on nonlinear manifolds due to orthonormality constraints. While the present work is formulated in a Lagrangian framework, it can be viewed as a precursor to a fully Riemannian treatment of the problem. A consistent definition of the energy gradient in the Riemannian sense has only recently been introduced for single-determinant methods [6], and its extension to the multiconfigurational setting is an important direction for future work. In the present work, however, we employ tools that are more familiar to specialists in quantum chemistry, namely, Lagrange’s multiplier theorem for the constrained optimization and the Hilbert space of square integrable functions.
@@
Our approach allows one to perform MCSCF calculations of both ground and excited states within the framework of MRA. Here, we note that this adaptive discretization technique is particularly useful in computational chemistry, where it effectively handles the singular Coulomb interaction between an electron and the nucleus, as well as the resulting cusp in orbitals [7, 9, 11, 23]. Moreover, its success in chemistry applications is largely due to the approximation of the resolvent:
(1.1) $H(\mu)=\left(-\Delta+\mu^{2}\right)^{-1}\approx\sum_{j}c_{j}e^{t_{j}\Delta}$
@@
based on the heat semigroup [8]. Multiresolution quantum chemistry is rapidly evolving, and we refer readers to a comprehensive review for a broader perspective [3].
We employ the Lagrange’s multiplier theorem for constrained minimization of the MCSCF energy. Although this approach is less common in quantum chemistry – since stationary points of the Lagrangian $\mathcal{L}$ are saddle points – it provides a natural framework for incorporating Green’s operators. As a result, the differential Newton system can be reformulated in an integral form, which is particularly well suited for implementation within the multiresolution framework. An alternative MRA approach to MCSCF problem can be seen in recent works [16, 22].
@@ 2. Notations
All the formulas are given in atomic units. Throughout the text we use the notation $\varphi,\varphi_{i},\delta\varphi_{i}$ for functions in $L^{2}\left(\mathbb{R}^{3}\right)$, which are referred to as orbitals. We focus on closed-shell two-electron systems. A Slater determinant associated with a spatial orbital $\varphi_{m}$ is denoted by $\left|m\overline{m}\right\rangle,$ indicating spin-up and spin-down electrons occupying the same spatial orbital. The one-electron Hamiltonian is defined as
@@
We employ the standard notation
$(i|h|j)=\int\varphi_{i}(x)h\varphi_{j}(x)\,dx$
for one-electron integrals and
$(ij|kl)=\int\frac{1}{|x-y|}\varphi_{i}(x)\varphi_{j}(x)\varphi_{k}(y)\varphi_{l}(y)\,dxdy$
for two-electron Coulomb integrals [21]. We frequently encounter convolution expressions of the form
$J(i\,\delta k)=\frac{1}{|x|}*(\varphi_{i}\,\delta\varphi_{k}),$
@@
which is a function in $L^{3*}\left(\mathbb{R}^{3}\right)$, known as weak $L^{3}$-space, by the Hardy–Littlewood–Sobolev lemma. In the computational chemistry literature, such expressions are naturally associated with Coulomb and exchange potentials, which motivates this notation.
Finally, we introduce the resolvent operator
(2.1) $R_{i}=\left(-\frac{c_{i}^{2}}{2}\Delta-\varepsilon_{ii}\right)^{-1}=\frac{2}{c_{i}^{2}}H\!\left(\sqrt{-\frac{2\varepsilon_{ii}}{c_{i}^{2}}}\right),$
where the orbital energies $\varepsilon_{ii}$ are assumed to be negative. This ensures that the orbital energies lie in the resolvent set of the kinetic energy.
In practical computations, this condition may be temporarily violated during early iterations. Such a situation indicates that the current iterate is far from convergence. In this case, one may enforce negativity by replacing $\varepsilon_{ii}$ with $\varepsilon_{ii}$. As will become clear below, this procedure is consistent with the standard level-shift strategy used to stabilize the convergence in Newton-type methods.
In general, an optimization problem can be recast as a stationary point equation
(2.2) $\mathcal{R}(w)=0$
with some $\mathcal{R}:W\to W$, where the space $W$ may contain functions, coefficients and Lagrange multipliers. Newton’s method can be described as follows. For a given iterate $w=w_{n}$ we solve the Newton’s equation
(2.3) $d\mathcal{R}(w)\delta w=-\mathcal{R}(w)$
with respect $\delta w$. This is a linear equation that we solve iteratively. Note that in a basis type method one could potentially invert the Jacobian $d\mathcal{R}(w)$, though for this the basis size should be prohibitively small. Then the update $\delta w$ gives the new iterate $w_{n+1}=w_{n}+\delta w$, which in turn defines the new Newton’s equation (2.3), now with $w=w_{n+1}$ and new unknown $\delta w$. In other words the complete iteration procedure consists of two loops. The innermost loop is solved with the help of fixed-point iterations. For this purpose we rewrite (2.3) in the so called self-consistent form
(2.4) $\delta\varphi=\mathbb{F}(\delta\varphi;w),$
where we separate the orbital update $\delta\varphi$ from the full update vector $\delta w$. In (2.4) the point $w$ is a fixed parameter, while $\delta\phi$ is unknown. One iterates $\delta\varphi_{n+1}=\mathbb{F}(\delta\varphi_{n};w)$ until self-consistency. It can be accelerated with direct inversion in the iterative subspace (DIIS), an extrapolation technique also known as Pulay mixing [17, 18]. Concrete examples of $\mathcal{R}$ in (2.3) and $\mathbb{F}$ in (2.4) are provided below in the text.
@@ 3. Two configurations
(3.1) $\frac{1}{4}\frac{\delta\mathcal{L}}{\delta\varphi_{1}}=c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}$
@@
(3.2) $c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}=\varepsilon_{11}\varphi_{1}+\varepsilon_{12}\varphi_{2}$
@@
$\sin 2\theta(2(1|h|1)+(11|11))+\sin 2\theta(2(2|h|2)+(22|22))+2\cos 2\theta(12|12)=0$
@@
acting in $L^{2}\left(\mathbb{R}^{3}\right)^{2}\times\mathbb{R}^{4}.$ We are interested in those roots of $\mathcal{R}$, points $w$ satisfying (2.2), which give the lowest possible energy. At each Newton step we have to deal with (2.3). The first two lines in the Newton system (2.3) have the form
(3.3) $d\mathcal{R}_{i}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{i}(w)\delta\varphi_{1}+\ldots+\frac{\partial\mathcal{R}_{i}}{\partial\varepsilon_{12}}(w)\delta\varepsilon_{12},\quad i=1,2,$
@@
$\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{11}}(w)\delta\varepsilon_{11}+\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{22}}(w)\delta\varepsilon_{22}+\frac{\partial\mathcal{R}_{2}}{\partial\varepsilon_{12}}(w)\delta\varepsilon_{12}=-\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}.$
@@
$\partial_{\varphi_{1}}\mathcal{R}_{3}(w)\delta\varphi_{1}=-4\sin 2\theta((\delta 1|h|1)+(\delta 11|11))+4\cos 2\theta(\delta 12|12),$
@@
and
$\frac{\partial\mathcal{R}_{3}}{\partial\theta}(w)\delta\theta=(-2\cos 2\theta(2(1|h|1)+(11|11))+2\cos 2\theta(2(2|h|2)+(22|22))-4\sin 2\theta(12|12))\delta\theta.$
Finally, we have
$d\mathcal{R}_{4}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{4}(w)\delta\varphi_{1}=2\int\varphi_{1}\delta\varphi_{1},$
@@
and
$d\mathcal{R}_{6}(w)\delta w=\partial_{\varphi_{1}}\mathcal{R}_{6}(w)\delta\varphi_{1}+\partial_{\varphi_{2}}\mathcal{R}_{6}(w)\delta\varphi_{2}=\int\delta\varphi_{1}\varphi_{2}+\int\varphi_{1}\delta\varphi_{2}.$
3.1. Self-consistent form
We reduce the full Newton system to a linear system in orbital updates $(\delta\varphi_{1},\delta\varphi_{2})$ by eliminating $\delta\theta$ and $\delta\varepsilon$. Firstly, $\delta\theta$ is defined as $\delta\theta=\Theta(\delta\varphi_{1},\delta\varphi_{2},w)$ by
$\delta\theta=-\left(\frac{\partial\mathcal{R}_{3}}{\partial\theta}(w)\right)^{-1}\left(\partial_{\varphi_{1}}\mathcal{R}_{3}(w)\delta\varphi_{1}+\partial_{\varphi_{2}}\mathcal{R}_{3}(w)\delta\varphi_{2}+\mathcal{R}_{3}(w)\right).$